### v1

# Sobolev inequality and its applications to nonlinear PDE on noncommutative Euclidean spaces

Abstract.
In this work, we study the Sobolev inequality on noncommutative Euclidean spaces. As a simple consequence, we obtain the Gagliardo–Nirenberg type inequality and as its application we show global well-posedness of nonlinear PDEs in the noncommutative Euclidean space. Moreover, we show that the logarithmic Sobolev inequality is equivalent to the Nash inequality for possibly different constants in this noncommutative setting by completing the list in noncommutative Varopoulos’s theorem in [37]. Finally, we present a direct application of the Nash inequality to compute the time decay for solutions of the heat equation in the noncommutative setting.
Key words and phrases:
Noncommutative Euclidean space, Sobolev inequality, Nash inequality, Logarithmic Sobolev inequality, Gagliardo–Nirenberg inequality, nonlinear damped wave equation, global well-posedness2020 Mathematics Subject Classification:
46L51, 47L25, 46E35, 42B15, 35L76, 35L05Contents
1. Introduction
The classical Sobolev inequality [31] in the Euclidean space states that, for any sufficiently smooth function with compact support, the inequality
holds, where the constant depends only on This inequality is widely applied in the study of partial differential equations (briefly PDEs) (e.g., [11]) and is interconnected with various other inequalities. The Sobolev inequality finds widespread applications in diverse areas of mathematics and physics. In the study of elliptic and parabolic partial differential equations, it serves as a fundamental tool for proving existence and uniqueness of solutions, as well as regularity properties of solutions. Moreover, in the context of variational problems and optimization, the Sobolev inequality plays a crucial role in establishing the compactness of minimizing sequences and the convergence of solutions. Extending this theory to noncommutative settings poses unique challenges and opportunities. In the noncommutative context, the traditional notion of differentiation must be reinterpreted, often relying on concepts from operator theory and functional analysis. In the framework of non-commutative geometry, the notion of Euclidean space undergoes a profound transformation. Instead of traditional coordinates that commute with each other, we deal with noncommuting coordinates, reflecting the noncommutative nature of the space. This departure from commutativity introduces intriguing mathematical structures, reminiscent of quantum mechanics and operator algebras. Noncommutative Euclidean space which is defined in terms of an arbitrary skew-symmetric real matrix represent deformations of Euclidean space as outlined by Rieffel [26]. This family of spaces stands among the earliest and most thoroughly investigated examples in noncommutative geometry, owing to its significance in the quantum mechanical phase space perspective. Various scholars, including Moyal [23] and Groenwald [13], have approached these spaces from diverse angles. The central concept involves deforming the algebra of smooth functions on by substituting the standard pointwise product with the twisted Moyal product. In the realm of noncommutative geometry, noncommutative Euclidean spaces represent notable instances of “noncompact” spaces [2, 9]. The Moyal product garners attention for its relevance to quantum phase space [15, Chapter 13], [12]. Furthermore, noncommutative Euclidean spaces play a crucial role in physics, especially in situations where spatial coordinates fail to commute [2].
In the literature, multiple equivalent constructions regarding noncommutative Euclidean space are documented. One approach to defining starts by establishing the von Neumann algebra as a twisted left-regular representation of on We will delve into this definition of a von Neumann algebra, which is generated by a -parameter strongly continuous unitary family satisfying the relation
where is inner product in
In recent times, there has been a significant body of research aimed at extending the techniques of the classical harmonic analysis on Euclidean spaces to noncommutative setting. This is because it is possible to define analogues of many tools of harmonic analysis, such as differential operators, Fourier transform, and function spaces. Recent progress in this theory even allow to study nonlinear PDEs [21]. For more details and recent results on this theory, we refer the reader to [8], [9], [20], [21], [16], [28], and references therein. At its core, the Sobolev inequality relates the smoothness of a function to the behavior of its derivatives. In particular, it quantifies the growth of the -norm of a function and its derivatives in terms of the -norm of the function itself, where and are related exponents depending on the dimension of the space and the order of differentiation.
There are several proofs of the Sobolev inequality in One of them is by using the heat kernel estimate [39]. Indeed, N. Varopoulos obtained an equivalence between a heat kernel estimate and the Sobolev inequality in an abstract settings in 1985 [38]. He proved that the Sobolev inequality is equivalent to the heat kernel estimate as well as the Nash, and Moser inequalities, for possibly different constants. On the other hand, the Nash inequality is also equivalent to the logarithmic Sobolev inequality. The Sobolev inequality was proved in [37, Theorem 3.11 and Corollary 3.19] by using the heat kernel estimate. Moreover, the author obtained the noncommutative analogue of the Varopoulos’s theorem even a more general case which includes noncommutative Euclidean spaces. On the other hand, there is another way to prove the Sobolev inequality which hinges upon the Young inequality for weak type spaces and the Hardy-Littlwood-Sobolev inequality. Recently the Sobolev inequality was also studied in the quantum phase space in [18] by this way. We think that the quantum phase space is equivalent to the particular case of our noncommutative Euclidean space when In this work, we study general case which covers some results in [18]. Therefore, we first focused to obtain the Hardy-Littlwood-Sobolev and Young inequality for weak type spaces in the noncommutative Euclidean space. Our proof is based on the Young inequality in [21] and the interpolation in the classical Lorentz spaces as well as a proper extension of the definition of convolutions in the class of distributions. Moreover, we prove that the equivalence between the Nash inequality the logarithmic Sobolev inequality remains true even in the noncommutative Euclidean space by completing the list of the noncommutative Varopoulos’s theorem in [37, Theorem 4.30]. As a consequence of the noncommutative Sobolev and Hölder inequalities we obtain a version of the Gagliardo–Nirenberg inequality which allows us to show its global well-posedness of nonlinear damped wave equations for the sub-Laplacian in the noncommutative Euclidean space. Nonlinear PDEs in the noncommutative Euclidean space were studied very recently in [21], where the author obtained local (some global) well-posedness in time of Allen-Cahn, Schrödinger and incompressible Navier-Stokes equations. The main difference in this work is that we study damped wave equations for the sub-Laplacian and show its global in time well posedness. Moreover, our approach is completely different from that of [21]. However, our approach is also applicable at least to obtain a global in time well posedness of the nonlinear Schrödinger equation which was studied in [21]. In general, there are some technical difficulties in noncommutative setting including defining even nonlinear operator functions. The other difficulties in the noncommutative Euclidean spaces were explained in [21]. The main tool to show that is the Gagliardo–Nirenberg type inequality and proper Banach fixed point theorem. The idea comes from the paper [29], where the authors studied similar problems in the context of Heisenberg and graded Lie groups. At the end, as in the classical case, we show a direct application of the Nash inequality to compute the time decay for solutions of the heat equation in the noncommutative setting.
2. Preliminaries
2.1. Noncommutative (NC) Euclidean space
For a thorough examination of the noncommutative (NC) Euclidean space and additional insights, we recommend recent works [8], [9], [16], [20], [21], and [28].
Let represent a Hilbert space, with denoting the algebra consisting of all bounded linear operators that act on . In the usual context, we denote by () the -spaces of pointwise almost-everywhere equivalence classes of -integrable functions, while denotes the space of essentially bounded functions on the Euclidean space .
For , the Lorentz space refers to the space of complex-valued measurable functions on satisfying the finiteness of the quasinorm defined as:
where represents the distribution of the function . For a more comprehensive understanding of these spaces, we refer the reader to [14].
Let us suppose we have an integer , and we choose an antisymmetric -valued matrix
Definition 2.1.
Define (or as the von Neumann algebra generated by the -parameter strongly continuous unitary family satisfying the relations
where means the usual inner product in
The relation mentioned above is known as the Weyl representation of the canonical commutation relation. While it is viable to define in an abstract operator-theoretic manner as outlined in [9], an alternative approach involves defining the algebra through a specific set of operators defined on the Hilbert space .
Definition 2.2.
[21, Definition 2.1] Let denote the operator on for , defined by:
It can be demonstrated that the family is strongly continuous and satisfies the relation (2.2). Subsequently, the von Neumann algebra is defined as the weak operator topology closed subalgebra of generated by the family , and is called a non-commutative (or quantum) Euclidean space.
It should be noted that when , the definitions provided above reduce to characterizing as the algebra of bounded pointwise multipliers on .
Generally, the algebraic nature of the noncommutative Euclidean space depends on the dimension of the kernel of In the case up to an orthogonal conjugation may be given as
for some constant In this case, is -isomorphic to and this -isomorphism can be written as
where and is the differentiation. If then an arbitrary antisymmetric real matrix can be expressed (up to orthogonal conjugation) as a direct sum of a zero matrix and matrices of the form (2.3), ultimately leading to the -isomorphism
where is the von Neumann tensor product [20]. In particular, if then (2.4) reduces to
It is worth noting that these formulas make sense due to the fact that the rank of an antisymmetric matrix is always even.
2.2. Noncommutative integration
Given we define the operator by the formula
This integral converges absolutely in the Bochner sense, yielding a bounded linear operator such that (see [20, Lemma 2.3]). Let us use to represent the classical Schwartz space on For any we define the Fourier transform as:
The noncommutative Schwartz space is precisely the set of elements of that can be expressed as for some belonging to the classical Schwartz space In other words:
We equip with a topology induced by the canonical Fréchet topology on via the map The topological dual of is denoted as Moreover, the injectivity of the mapping is established [20, Subsection 2.2.3], [21], and for it can be extended to distributions
For any , we define the functional with the formula:
2.3. Noncommutative and spaces
Given the definitions outlined in the earlier sections, emerges as a semifinite von Neumann algebra, with serving as its trace. Thus, the pair is characterized as a noncommutative measure space. We can define the -norm on this space for any using the Borel functional calculus and the following expression:
where The space is the completion of the set under the norm and this completion is denoted by The elements of are -measurable operators, like in the commutative case. These are linear densely defined closed (possibly unbounded) affiliated with operators such that for some Here, denotes the spectral projection corresponding to the interval Let represent the collection of all -measurable operators. We assume . The distribution function of is defined by
For we define the generalised singular value function as follows:
The function is decreasing and right-continuous. For further discussion on generalized singular value functions, we direct the reader to [7, 19]. The norm of can also be represented using the generalized singular value function, as outlined in (see [19, Example 2.4.2, p. 53]):
The equality for was established in [19, Lemma 2.3.12. (b), p. 50].
The space is a -algebra, which can be endowed with a topological -algebra structure in the following manner. We consider the set
Then, the family constitutes a system of neighbourhoods at 0, resulting in becoming a metrizable topological -algebra. The convergence induced by this topology is called the convergence in measure [24]. Next, we establish the noncommutative Lorentz space linked with the noncommutative Euclidean space.
For we introduce the non-commutative Lorentz space as follows:
2.4. Differential calculus on
Let us recall the differential structure on as detailed in (see, [21, Subsection 2, p. 10]).
The differential structure relies on the group of translations where is presented as the unique -automorphism of that operates on as follows:
where denotes the standard inner product in
Expressed in terms of the map we observe:
for all
Alternatively, for we can introduce as the conjugation of by the unitary operator representing translation by on
Definition 2.3.
(see, [21, Definition 2.9]) An element is considered smooth if, for all the function is smooth.
The partial derivatives where are defined on smooth elements as follows:
Using (2.6) and (2.11), we can readily confirm that for the following holds:
We introduce the notation for a multi-index as follows:
and denote the gradient associated with as defined as the operator
Furthermore, we define the Laplace operator as
Through the concept of duality, we can further extend the derivatives to operators on
In a manner of the classical setting, let us denote the pairing between from and from as We embed the space into via:
For a multi-index and the distribution is defined as:
Moreover,
2.5. Fourier transform on NC Euclidean spaces
Definition 2.4.
For any we define the Fourier transform of as the map by the formula
Moreover, we have Plancherel’s (Parseval’s) identity [20]
Lemma 2.5.
Here, is the spectrum of
Remark 2.6.
The inequality (2.18) is of course reversed if we employ an operator concave function instead of an operator convex function.
2.6. Convolution on
Definition 2.7.
Let and For we define
where the integral is understood in the sense of a - valued Bochner integral when and as a weak* integral when
The following remark can be found from [21].
Remark 2.8.
The convolution (2.19) preserves the positivity: if and then
Definition 2.9.
For any define
The following proposition shows that this is, indeed, well defined and belongs to the class of Schwartz functions.
Proposition 2.10.
If then as a function belongs to In other words,
Proof.
Let Then we can write and for some and in respectively. Moreover, we obtain the following equality (see, [20, formula 2.13])
where the -convolution as
It follows from (2.12) and (2.20) that
Therefore,
where
Since it follows that belongs to , thereby completing the proof. ∎
Definition 2.11.
Let and Then we define a convolution in by
For example, if is the Dirac mass, then
Hence,
The following noncommutative analogue of the Varopoulos’s theorem in the noncommutative Euclidean space can be inferred from [37].
Theorem 2.12.
[37, Theorem 3.11 and Corollary 3.19] Let Then the following inequalities are equivalent,
with possibly different values of the constants
3. Hardy–Littlewood–Sobolev inequality
In this section, we prove a quantum analogue of the Hardy–Littlewood–Sobolev inequality. First, we prove a noncommutative version of [14, Theorem 1.2.13, p. 23] for weak- spaces.
Theorem 3.1.
(Young inequality for weak type spaces) Let and satisfying
Then there exists a constant such that
for all and
Proof.
Let be a positive real number to be chosen later. Set and In view of [14, Exercise 1.1.10(a), p. 23] we obtain
and
Here, is the the classical distribution function of (see, [14, Definition 1.1.1. p. 2]). Then by [14, formulas (1.2.20) and (1.2.20). p. 24], we obtain
for and
for
Since we find Now, let us choose and Hence, it follows from the noncommutative Young inequality [21, Theorem 3.2, p. 16] and (3.4) that
If then we choose such that the right-hand side of (3.6) is equal to that is
and if for In this case, since we have that In this instance, we utilized this fact, if and then it is clear that if and only if and (see the proof of [6, Lemma 3.2.3, p. 125]).
Let Then it follows from the noncommutative Young inequality [21, Theorem 3.2, p. 16] with and (3.5) that
Therefore, for the value of chosen, applying (3.7) and the noncommutative Chebyshev inequality [30, Lemma 4.7, p. 4094], we have
In other words, we have
Taking supremum over from both sides of the previous inequality we obtain the desired result. ∎
Next, we establish a noncommutative version of [14, Theorem 1.4.25. p. 73] for weak- spaces.
Theorem 3.2.
(Hardy–Littlewood–Sobolev inequality). Let satisfying
Then, there exists such that
for and
Proof.
Lemma 3.3.
Let Then for any and we have
where is the classical Laplacian.
Proof.
Let and For set
Hence, it follows from (2.12), (3.10), and (2.20) that
and
Therefore,
and
Combining (2.24), (3.10), (2.23), and (3.14), we obtain
In other words, we have
This is the second equality. Let us now show that
Indeed, by using (3.10), (2.12), and (2.13), we find
Therefore,
Hence, we have
for which gives (3.17). Then the combination of (3.16) and (3.17) gives the desired result. ∎
Remark 3.4.
Define the function on by
with constants
and
respectively, (see, [14, Theorem 2.4.6. p. 138]). In the case it is well-known that this function gives the fundamental solution of the Laplacian (see, [14, Proposition 2.4.4. p. 135]), that is,
The reflection is defined as follows (see, [14, Definition 2.3.11])
Next we give an extension of the property of the Fourier transform to tempered distribution (see, [14, Proposition 2.3.22.]):
It is follows from (2.7), (2.24) (3.14), (3.20) and (3.21) that
For using (3.15) we find
Thus,
Theorem 3.5.
(Sobolev type inequalities) Let and satisfy
There exist a constant such that for any we have
Proof.
As a particular case of inequality (3.24), when and such that we establish the following Sobolev inequality.
Corollary 3.6.
(Sobolev inequality). Let with Then there exist a constant such that for any we have
4. Equivalence of various inequalities
4.1. The equivalence of Nash inequality and logarithmic Sobolev inequality
In this subsection, we obtain the logarithmic Sobolev inequality. The logarithmic Sobolev inequality was obtained recently in [28] on noncommutative Euclidean spaces when and with a restriction for the parameter By using the Sobolev inequality we prove the inequality without these restrictions and complete the range of for
Theorem 4.1.
(Logarithmic Sobolev inequality). Let be such that Then for any such that we have the fractional logarithmic Sobolev inequality
where is a constant independent of and
Proof.
Next, we show that the Nash type inequality and the logarithmic Sobolev inequality are equivalent when
Theorem 4.2.
Let Then for nay such that the following inequalities are equivalent:
and
Proof.
Let us show that inequality (4.1) implies inequality (4.2). We assume that Applying the Jensen inequality for the convex function we obtain
Note that the map is a positive normalized linear map. Moreover, we know that the logarithmic function is operator concave on (see [25, Example 1.7]). Therefore, as we have it follows from Jensen operator inequality (2.18) and Remark 2.6 that we have
By simple properties of the logarithmic function and inequality (4.1), we can write
Combining (4.3), (4.4) and (4.5), we infer that
Thus, by taking both sides of the previous inequality to the power of we obtain (4.2).
Conversely, let us consider the non-commutative Hölder inequality (see [28, Lemma 6.3]):
with such that Hence, by taking the logarithm from both sides of the inequality (4.6) we obtain
This inequality becomes an equality when Moreover, we can rewrite the inequality (4.6) as follows
For the Logarithmic Hölder inequality (see, [28, Lemma 6.3]) holds
Since
and
differentiating the function with respect to and we have
From the fact and the estimate (4.9) we get
In other words,
If and then we can rewrite the inequality (4.10) as follows
Next, by the Nash type inequality (4.2), we get
Theorem 4.3.
Let Then for any such that the following inequalities are equivalent:
and
5. Applications to Nonlinear PDEs
5.1. A Gagliardo–Nirenberg type inequality
We will now apply the Sobolev inequality to prove a Gagliardo–Nirenberg type inequality on the noncommutative Euclidean space. Further, we apply it to well-posedness of nonlinear PDEs. Then the following Gagliardo–Nirenberg type inequality holds.
Theorem 5.1.
Let and Assume that
Then we have the following Gagliardo–Nirenberg type inequality
where and
Proof.
As a special case of Theorem 5.1 when we obtain the following inequality which plays a key role for our further investigations.
Corollary 5.2.
Let and Then we have the following inequality
for
5.2. Linear damped wave equation on noncommutative Euclidean spaces
Now, we consider the linear equation
with the Cauchy data
and
where the damping is determined by and the mass
Set and where
For the solution of the initial-value problems (5.12) for each we obtain explicit formulas
and
for each fixed From this we have the estimates
and
respectively. Hence, repeating the argument in [29, pp. 5221-5223], there is a positive constant such that in all cases, we have
The Sobolev space with associated to operator is defined as
Proposition 5.3.
Let and assume that and Then there exists a positive constant such that
holds for all Moreover, for any we have
for all
Proof.
By the estimate (5.13) we have
for all Moreover, for any and we have
thereby completing the proof. ∎
5.3. Semilinear damped wave equations on noncommutative Euclidean spaces
Let be a function satisfying the condition
for some and self-adjoint operators
The absolute value function is Lipschitz continuous on all -spaces for as demonstrated in the self-adjoint case in [5]. For the case of normal operators, refer to the paper by [17]. This result enables the application of the absolute value function to your operators and allowing you to focus exclusively on the case of positive operators.
The class of functions satisfying condition (5.18) is not empty. For example, the function satisfies this condition for even for normal operators. Indeed, it is clear that and by the noncommutative Hölder inequality (see, [33, Theorem 1]), we obtain
for any normal operators Repeating this argument inductively we can also prove that the functions such as satisfy condition (5.18) for positive operators and for even integer Now, we consider a semilinear wave equation for the operator on the noncommutative Euclidean space with the nonlinearity satisfying condition (5.18).
Let us study
with the damping term determined by and with the mass
We now formulate our main result in the noncommutative Euclidean space.
Theorem 5.4.
Let the Cauchy data and satisfy
Then, there exists a small constant such that the Cauchy problem (5.22) has a unique global solution for all
Moreover, there exists a number such that
Proof.
Let us consider a closed subset of the space defined by
with
where will be defined later. Define the mapping on by the formula
where is the solution of the problem (5.6)-(5.8), and is the solution of the following linear problem:
Moreover, we can present the operator as follows
where
We assume for a moment that
and
Hence, it follows from (5.28) and (5.29) that is a contraction mapping on The Banach fixed point theorem then implies that has a unique fixed point in It means that there exists a unique global solution of the equation
which also gives the solution to (5.22). Therefore, we have to prove inequalities (5.28) and (5.29).
By the condition (5.18) and the Gagliardo–Nirenberg inequality (5.5), and applying the following Young inequality
and we obtain
Since and it follows that
By putting in (5.30), and using that we also have
Now, we consider the operator By the Fourier transform (2.16) and (5.27), we obtain
for all
Hence, it follows from Proposition 5.3 and the Plancherel (Parseval) identity (2.17) and (5.32) that
for
Consequently, using (5.14), (5.15), (5.25) and (5.35), we obtain
for some and Moreover, in the similar way, we can estimate
for some Take and choose with sufficiently small so that
Hence, by using formulas (5.36)–(5.38) we obtain the estimates (5.28) and (5.29). This allows us to apply the fixed point theorem to establish the existence of solutions. Moreover, the estimate (5.24) follows from (5.33), thereby completing the proof. ∎
5.4. Application of the Nash inequality to the heat equation
In this subsection we present a direct application of the Nash inequality to compute the decay rate for the heat equation with as in (2.14).
Theorem 5.5.
Proof.
The solution of equation (5.39) is written as
Since and is a positive operator, it follows from Remark 2.8 that is a positive operator for any Hence, by (2.19) and (5.40) we have
which implies that
Therefore, we have
Set Then applying the Nash inequality (4.2) with , and by (5.42) and (5.41) we obtain
Integrating with respect to we obtain the following estimate
which completes the proof. ∎
6. Acknowledgements
Authors would like to thank to Dr. Edward McDonald for helpful discussions on non-commutative Euclidean spaces. Indeed, he offered the new definition of the convolution in Definition 2.9. The work was partially supported by the grant from the Ministry of Science and Higher Education of the Republic of Kazakhstan (Grant No. AP23487088). The authors were partially supported by Odysseus and Methusalem grants (01M01021 (BOF Methusalem) and 3G0H9418 (FWO Odysseus)) from Ghent Analysis and PDE center at Ghent University. The first author was also supported by the EPSRC grant EP/V005529/1.

### v5

@@
 On the other hand, the Nash inequality is also equivalent to the logarithmic Sobolev inequality.
-The Sobolev inequality was proved in [37, Theorem 3.11 and Corollary 3.19] by using the heat kernel estimate.
+The Sobolev inequality was proved in [38, Theorem 3.11 and Corollary 3.19] by using the heat kernel estimate.
 Moreover, the author obtained the noncommutative analogue of the Varopoulos’s theorem even a more general case which includes noncommutative Euclidean spaces.
@@
 Our proof is based on the Young inequality in [21] and the interpolation in the classical Lorentz spaces as well as a proper extension of the definition of convolutions in the class of distributions.
-Moreover, we prove that the equivalence between the Nash inequality the logarithmic Sobolev inequality remains true even in the noncommutative Euclidean space by completing the list of the noncommutative Varopoulos’s theorem in [37, Theorem 4.30].
+Moreover, we prove that the equivalence between the Nash inequality the logarithmic Sobolev inequality remains true even in the noncommutative Euclidean space by completing the list of the noncommutative Varopoulos’s theorem in [38, Theorem 4.30] (see also [34] for the equivalence between -ultracontractivity and logarithmic Sobolev inequality for some in noncommutative setting).
 As a consequence of the noncommutative Sobolev and Hölder inequalities we obtain a version of the Gagliardo–Nirenberg inequality which allows us to show its global well-posedness of nonlinear damped wave equations for the sub-Laplacian in the noncommutative Euclidean space.
@@
 Theorem 2.12.
-[37, Theorem 3.11 and Corollary 3.19] Let Then the following inequalities are equivalent,
+[38, Theorem 3.11 and Corollary 3.19] Let Then the following inequalities are equivalent,
 with possibly different values of the constants
@@
 4. Equivalence of various inequalities
-4.1. The equivalence of Nash inequality and logarithmic Sobolev inequality
+4.1. Equivalence of the Nash and the logarithmic Sobolev inequalities
 In this subsection, we obtain the logarithmic Sobolev inequality.
@@
 Theorem 4.1.
-(Logarithmic Sobolev inequality). Let be such that Then for any such that we have the fractional logarithmic Sobolev inequality
+(Logarithmic Sobolev inequality). Let be such that Then for any we have the fractional logarithmic Sobolev inequality
 where is a constant independent of and
@@
 Theorem 4.2.
-Let Then for nay such that the following inequalities are equivalent:
+Let Then for any the following inequalities are equivalent:
 and
@@
 and
-differentiating the function with respect to and we have
+differentiating the function with respect to (see [34, Lemma 4.2]), and we have
 From the fact and the estimate (4.9) we get
@@
 Theorem 4.3.
-Let Then for any such that the following inequalities are equivalent:
+Let Then for any the following inequalities are equivalent:
 and
@@
 respectively. Hence, repeating the argument in [29, pp. 5221-5223], there is a positive constant such that in all cases, we have
-The Sobolev space with associated to operator is defined as
+The Sobolev space with associated to the operator is defined as
 Proposition 5.3.
@@
 The absolute value function is Lipschitz continuous on all -spaces for as demonstrated in the self-adjoint case in [5].
-For the case of normal operators, refer to the paper by [17].
-This result enables the application of the absolute value function to your operators and allowing you to focus exclusively on the case of positive operators.
+For the case of normal operators, we refer the reader to the paper [17].
+These results enable the application of the absolute value function to operators and allowing us to focus exclusively on the case of positive operators.
 The class of functions satisfying condition (5.18) is not empty.
@@
 By the condition (5.18) and the Gagliardo–Nirenberg inequality (5.5), and applying the following Young inequality
-and we obtain
+we obtain
 Since and it follows that
@@
 which completes the proof. ∎
+5.5. Conflict of interest
+We can conceive of no conflict of interest in the publication of this paper. The work has not been published previously and it has not been submitted for publication elsewhere.
+5.6. Data availability
+No new data were created or analysed during this study. Data sharing is not applicable to this article
 6. Acknowledgements